### v1

@@ Abstract
Communication overhead is a key challenge in distributed deep learning, especially on slower Ethernet interconnects, and given current hardware trends, communication is likely to become a major bottleneck. While gradient compression techniques have been explored for SGD and Adam, the Lion optimizer has the distinct advantage that its update vectors are the output of a sign operation, enabling straightforward quantization. However, simply compressing updates for communication and using techniques like majority voting fails to lead to end-to-end speedups due to inefficient communication algorithms and reduced convergence. We analyze three factors critical to distributed learning with Lion: optimizing communication methods, identifying effective quantization methods, and assessing the necessity of momentum synchronization. Our findings show that quantization techniques adapted to Lion and selective momentum synchronization can significantly reduce communication costs while maintaining convergence. We combine these into Lion Cub, which enables up to 5x speedups in end-to-end training compared to Lion. This highlights Lion’s potential as a communication-efficient solution for distributed training.
footnotetext: 1Institute of Science Tokyo 2Lawrence Livermore National Laboratory. Correspondence to: Anonymous Author <anon.email@domain.com>.
Preliminary work.
@@ 1 Introduction
Many algorithmic approaches have been developed to reduce communication volume, typically via gradient compression (Tang et al., 2020) in the context of SGD. These include sign-based algorithms (e.g., Seide et al., 2014; Bernstein et al., 2018; 2019), which communicate only the gradient sign; sparsification (e.g., Ström, 2015; Dryden et al., 2016; Lin et al., 2018; Renggli et al., 2019); and other methods with randomized compression operators (e.g., Alistarh et al., 2017; Wen et al., 2017; Wang et al., 2018). Similar ideas have been applied to more complex optimizers, such as Adam or LAMB (Tang et al., 2021; Li et al., 2022; Lu et al., 2023). However, these methods often slow down convergence due to information loss during compression, although recent papers have mitigated this effect.
@@ 2 Background and Related Work
SignSGD
SignSGD is an optimization method that applies the $\operatorname{sign}$ operator to gradients, resulting in three possible values: +1, 0, and -1 (Seide et al., 2014; Bernstein et al., 2018; 2019). This approach has gained attention due to its similarity to Adam, particularly in terms of the normalized update (Kunstner et al., 2023), but also as a distributed learning method when combined with majority voting, as updates can be compressed (Bernstein et al., 2019). In majority voting, $\operatorname{sign}$ is first applied to the gradients, and then again to their sum:
$\theta_{t+1}=\theta_{t}-\eta_{t}\operatorname{sign}\left[\sum_{i=1}^{N}\operatorname{sign}\nabla_{\theta}f_{i}(\theta_{t-1})\right].$
Note that $\operatorname{sign}$ is unbiased, which is why a stochastic $\operatorname{sign}$ is sometimes employed (Safaryan & Richtárik, 2021). Typically, discussions of majority voting assume a parameter server model for communication, where the signs computed on each worker are sent to a central server which computes the global sign and returns the result. As we will discuss, while straightforward, this approach does not scale.
Algorithm 1 Distributed Lion with Quantization
1:  given $\beta_{1},\beta_{2},\lambda,\eta,f,Q$
2:  initialize $\theta_{0},m_{0}\leftarrow 0$
3:  while $\theta_{t}$ not converged do
4:     $c_{i,t}\leftarrow\beta_{1}m_{i,t-1}+(1-\beta_{1})\nabla_{\theta}f_{i}(\theta_{t-1})$
5:     $c_{t}^{*}\leftarrow\sum_{i=1}^{N}Q(c_{i,t})$ sync update vector
6:     $\theta_{t}\leftarrow\theta_{t-1}-\eta_{t}\cdot\text{sign}(c_{t}^{*})+\lambda\theta_{t-1}$
7:     $m_{i,t}\leftarrow\beta_{2}m_{i,t-1}+(1-\beta_{2})\nabla_{\theta}f_{i}(\theta_{t-1})$
8:  end while
@@
where $\eta_{t}$ is the learning rate, $\beta_{1}$ and $\beta_{2}$ are hyperparameters (recommended defaults $0.9$ and $0.99$, resp.), and $m_{t}$ the momentum. When $\beta_{1}=\beta_{2}$, Lion matches sign descent. However, when $\beta_{2}>\beta_{1}$, the importance of the current gradient $\nabla f_{t}$ is increased compared to sign descent with momentum (Liu et al., 2024). Recently, Liu et al. (2024) proposed Distributed Lion (D-Lion, Algorithm 1), which combines Lion with majority voting, and demonstrated its convergence both theoretically and experimentally.
@@
Quantization maps signals from a larger set to a smaller one to achieve a more compact representation. A simple example is quantization to 1 bit using $x\mapsto x\geq 0$. When quantizing to $n$ bits, it is essential to ensure the full range of the representation is used. This is typically achieved by normalizing and rescaling the input data by the absoute maximum of the input elements, e.g.:
$Q_{\infty}(x)=\operatorname{sround}\left(\frac{2^{n-1}-1}{\|x\|_{\infty}}x\right)$
where $\|\cdot\|_{\infty}$ is the maximum norm and $\operatorname{sround}$ is a stochastic rounding operator (Alistarh et al., 2017).
This simple quantization method performs well with a uniform distribution, but frequently fails when applied to data following other distributions. Alternative methods have been proposed to handle non-uniform data distributions (e.g., Dettmers et al., 2022; 2023). Empirical evidence shows that gradients in deep learning frequently follow a Laplace or lognormal distribution, leading to methods specifically tailored to these distributions (e.g., Yu et al., 2020; Chmiel et al., 2021) but which are computationally intensive.
@@ 3 Communication Design
We first discuss the overall design and analysis of the communication primitives for Lion, and then on the implementation of our 1-bit allreduce in particular.
3.1 On communication for majority voting
@@
1-bit compressed allreduce $(P-1+\log_{2}(P))\alpha$ $(1+\frac{P-1}{P})N\beta$
@@
Figure 2: Communication costs with our performance model. We illustrate the communication costs from Table 1 for different scenarios of relative latency and bandwidth. In high-latency or large-scale settings, a direct is faster; conversely, in bandwidth-bound regimes, the 1-bit allreduce is faster.
@@
Lastly, we can adapt the 1-bit compressed allreduce of Tang et al. (2021) to support majority voting. The update vector is first represented to 1 bit (again forcing 0 to be either -1 or 1). This is then split into $P$ equal chunks, and the $i$th chunk is sent to the $i$th worker via an all-to-all communication. Each worker locally sums the contributions, takes the sign, and recompresses the result to 1 bit. Finally, an allgather is performed to disseminate the complete update to every worker. (See Fig. 1 for an illustration.) We assume the all-to-all is implemented as a pairwise exchange with cost $N\beta$, and the allgather has cost $\frac{P-1}{P}N\beta$, yielding the lowest bandwidth cost, although care must be taken to efficiently implement the intermediate packing and unpacking.
@@
We note that the 1-bit compressed allreduce logically uses an all-to-all to implement a reduce-scatter operation. However, this operation cannot be directly implemented with the standard reduce-scatter in NCCL or MPI due to the need to change bitwidths at intermediate reduction stages. A custom implementation doing this internally may yield further performance improvements; we leave this to future work.
3.2 The impact of 1-bit allreduces

### v2

@@ Abstract
Communication overhead is a key challenge in distributed deep learning, especially on lower-bandwidth interconnects, and given current hardware trends, communication is likely to become a major bottleneck. While gradient compression techniques have been explored for SGD and Adam, the Lion optimizer has the distinct advantage that its update vectors are the output of a sign operation, enabling straightforward quantization. However, simply compressing updates for communication and using techniques like majority voting fails to lead to end-to-end speedups due to inefficient communication algorithms and reduced convergence. We analyze three factors critical to distributed learning with Lion: optimizing communication methods, identifying effective quantization methods, and assessing the necessity of momentum synchronization. Our findings show that quantization techniques adapted to Lion and selective momentum synchronization can reduce communication costs while maintaining convergence. We combine these into Lion Cub, which enables speedups in training compared to Lion especially with low-bandwidth environments. This highlights Lion’s potential as a communication-efficient solution for distributed training.
Machine Learning, ICML
@@ 1 Introduction
Many algorithmic approaches have been developed to reduce communication volume, typically via gradient compression (Tang et al., 2020) in the context of SGD. These include sign-based algorithms (e.g., Seide et al., 2014; Bernstein et al., 2018, 2019) that communicate only gradient sign; sparsification (e.g., Ström, 2015; Dryden et al., 2016; Lin et al., 2018; Renggli et al., 2019); and randomized compression operators (e.g., Alistarh et al., 2017; Wen et al., 2017; Wang et al., 2018). Similar ideas have been applied to more complex optimizers, such as Adam or LAMB (Tang et al., 2021; Li et al., 2022; Lu et al., 2023). However, these methods often slow down convergence due to information loss during compression, although recent papers have mitigated this effect.
@@ 2 Background and Related Work
where $\eta_{t}$ is the learning rate, $\beta_{1}$ and $\beta_{2}$ are hyperparameters , and $m_{t}$ the momentum. When $\beta_{1}=\beta_{2}$, Lion matches sign descent. However, when $\beta_{2}>\beta_{1}$, the importance of the current gradient $\nabla f_{t}$ is increased compared to sign descent with momentum (Liu et al., 2024). Recently, Liu et al. (2024) proposed Distributed Lion, which combines Lion with majority voting, and demonstrated its convergence both theoretically and experimentally.
@@
When quantizing to $n$ bits, it is essential to ensure the full range of the representation is used. This is typically achieved by normalizing and rescaling the input data by the absolute maximum of the input elements, e.g.: $Q_{\infty}(x)=\operatorname{sround}\left(\frac{2^{n-1}-1}{\|x\|_{\infty}}x\right)$ where $\|\cdot\|_{\infty}$ is the maximum norm and $\operatorname{sround}$ is a stochastic rounding operator (Alistarh et al., 2017). This simple quantization method performs well with a uniform distribution, but frequently fails when applied to data following other distributions (Dettmers et al., 2022, 2023). Empirical evidence shows that gradients in deep learning frequently follow a Laplace or lognormal distribution, leading to methods specifically tailored to these distributions (e.g., Yu et al., 2020; Chmiel et al., 2021) but which are computationally intensive.
@@ 3 Communication Design
1-bit compressed allreduce $(P-1+\log_{2}(P))\,\alpha$ $\bigl{(}1+\frac{P-1}{P}\bigr{)}N\,\beta$
@@
Figure 2: Communication costs with our performance model. We illustrate the communication costs from Table 1 for different cases of relative latency and bandwidth.
@@
Lastly, we can adapt the 1-bit compressed allreduce of Tang et al. (2021) to support majority voting. The update vector is first represented to 1 bit (again forcing 0 to be either -1 or 1). This is then split into $P$ equal chunks, and the $i$th chunk is sent to the $i$th worker via an all-to-all communication. Each worker locally sums the contributions, takes the sign, and recompresses the result to 1 bit. Finally, an allgather is performed to disseminate the complete update to every worker. (See Figure 1 for an illustration.) We assume the all-to-all is implemented as a pairwise exchange with cost $N\beta$, and the allgather has cost $\frac{P-1}{P}N\beta$, yielding the lowest bandwidth cost, although care must be taken to efficiently implement the intermediate packing and unpacking.
@@
4 Quantization Design
4.1 The distribution of update vectors
Figure 3: The distribution of the update vector is not uniform. The distribution represents the values of the update vector before quantization ($\operatorname{sign}$). Left: The values do not follow a uniform distribution; rather, they approximate a Laplace distribution (not a Gaussian distribution). Right: The norm of the update vector is heavy-tailed, with the maximum significantly exceeding the mean.
4.2 L1 Quantization algorithm
Table 2: Validation loss on the Fineweb dataset with different quantization function.
Optim 32 bit 1 bit 8 bit ($L_{0}$) 8 bit ($L_{\infty}$) 8 bit ($L_{1}$)
Loss 2.94 3.22 3.07 4.64 2.97
We consider a new type of quantization function that normalizes by the $L_{p}$ norm instead of the $L_{\infty}$ norm, making it robust against outliers:
[…]
where $M_{p}(x)=\Bigl{(}\tfrac{1}{d}\sum_{j=1}^{d}|x_{j}|^{p}\Bigr{)}^{\!1/p}$ and $\operatorname{clamp}(\cdot,\alpha)$ is to truncate the values into $[-\alpha,\alpha]$. We call this quantization as $L_{p}$ quantization where $p=\infty$ corresponds to the standard quantization methods. As shown in Table 2, when training GPT on the FineWeb dataset, Lion converges poorly under $Q_{\infty}$ and 1-bit quantization, whereas $L_{0}$ and $L_{1}$ quantizations converge more effectively. Based on these findings, we will adopt $Q_{1}$ which uses the $L_{1}$ norm and is henceforth referred to as “L1 quantization” for all subsequent experiments. This quantization method is simple yet sufficiently effective for practical applications, as will be demonstrated later. Its simplicity also results in low computational overhead.
Figure 4 shows the accuracy of quantizing Lion updates with different methods in terms of whether the signs of the updates match or flip (introducing error). Standard $Q_{\infty}$ stochastic quantization clips most values to zero, resulting in a low match and flip rate. One approach to address this (“w/o zero”) is to stop rounding to zero and instead quantize sufficiently small values to $+1$ or $-1$. While this improves the match rate, it does not match distributed Lion. We also consider transforming values to a logarithmic scale before quantizing them, then reverting, which also does not match distributed Lion. In contrast, our L1 quantization significantly improves the sign match rate and reduces the flip rate compared to distributed Lion. We note that it may be surprising that distributed Lion is able to converge when its sign matching rate is 60–70%, only a little better than random. However, most of its errors are from returning zero, and consequently, its sign flip rate is around 10%. Further, the parameters that result in zeros often exhibit significant inter-worker variability, suggesting the need for updates in these cases is inherently low.
Figure 4: L1 quantization approximates Lion well. Top: Percent of update elements whose sign after quantization matches the standard Lion. Bottom: Percent of update elements whose sign after quantization reverses compared to standard Lion, changing the update direction. (This excludes elements mapped to 0.)
Figure 5: (Left) Momentum synchronization or L1 quantization is necessary when $\beta_{2}$ is small. Momentum is never synchronized for “None”; all other settings synchronize momentum every 10 iterations. When $\beta_{2}$ is large, gradients are stable across workers, and synchronization is not needed. When $\beta_{2}$ is small, synchronization is necessary for key layers. In addition, unlike standard distributed Lion, L1 quantization achieves convergence comparable to synchronized momentum, even without synchronization (“None”).
5 On the necessity of momentum synchronization
In standard Lion, while the final update vectors are 1-bit, the momentum is computed with full-precision gradients. In contrast, in distributed Lion (Liu et al., 2024), only the 1-bit updates are communicated, and the momentum term cannot be recovered, which can lead to momentum divergence between workers. Further, communicating the momentum would add overhead during training. We now investigate the extent to which momentum diverges and how to efficiently mitigate any divergence.
5.1 Layerwise variance of momentum
We can also see from Figure 5 that synchronizing the momentum of only the input and final layers is sufficient to achieve convergence similar to that of synchronizing the momentum across all layers at $\beta_{2}=0.95$. Figure 6 compares the variance in momentum between workers, showing, for each layer, the maximum standard deviation measured for the element with the highest deviation across workers. The standard deviation for the input embedding layer and the output layer is nearly 100 times larger than that of other layers. This indicates that the momentum in these layers varies significantly between workers, whereas there is almost no variation in the other layers. Given that the input and output layers are the layers where data directly enters, this result may be expected, considering that the data differs across workers. In vision models, it is only the input layer that exhibits such a large standard deviation (see Figure 16). We leave a more detailed analysis of this to future research.
5.2 Impact of quantization on momentum variance
These findings suggest two approaches for good convergence: 1-bit communication with momentum synchronization or L1 quantization with no momentum synchronization. The best choice depends on the specific setting. As a general guideline, when there are few workers, using L1 quantization is recommended. However, with a larger number of workers, the increased bitwidths required to correctly represent the majority vote limit potential speedups in training and quantization information. For example, with 125 workers, a uint8 can can only support 0/1 summation and it becomes necessary to synchronize momentum occasionally.
6 Lion Cub
6.1 Performance results
[…]
6.2 Training results
We now validate that Lion Cub preserves convergence and model quality. Figure 8 presents training curves for GPT (774M) trained on the OpenWebText dataset and Fineweb dataset with 8 workers on 2 nodes. Lion Cub (8-bit) maintains convergence comparable to the 32-bit Lion, while Lion Cub (4-bit) shows a decline in convergence. Consequently, Lion Cub (8-bit) with L1 quantization achieves faster real-time convergence. The convergence quality of Lion Cub (4-bit) can be recovered by synchronizing the momentum of the head layer every ten iterations, without significantly impacting training time. On the other hand, synchronizing momentum across all layers creates a communication bottleneck, which limits the speedup in real-time training performance. We trained on the Fineweb dataset with more aggressive hyperparameter settings than those used for the OpenWebText dataset. Under these conditions, LionCub(1-bit) failed to converge, forcing us to adopt LionCub(4-bit). Because this setup is highly sensitive to the numerical precision, compressing gradients from 32-bit to 16-bit (bfloat16) prior to communication (the 16-bit Lion) resulted in unstable training. Overall, Lion Cub, employing L1 quantization, provides substantial benefits in both training stability and communication speed.
We also evaluate Lion Cub in two additional contexts for vision tasks in Figure 9. The first task is training ResNet-18 on CIFAR100 for 100 epochs. Second, we fine-tune a DeiT-Tiny model, pretrained on ImageNet, on CIFAR100 for 20 epochs. In both cases, Lion Cub achieves comparable convergence to Lion in less overall time. Further, this is achieved even when not using L1 quantization or momentum synchronization. Hence, it is possible to lower the communication volume significantly to reduce overall training time in these cases.
Figure 8: Validation loss with GPT (774M) when training with 8 workers on 2 nodes in our Ethernet configuration. Lion Cub reduces training time compared to standard Lion and can maintain comparable convergence. Note in 4-bit Lion Cub, workers only send 1/-1 values, whereas in 8-bit Lion Cub, we use L1 quantization and send values in $[-15,15]$. In 4-bit, sync head Lion Cub, we synchronize the momentum of the final layer every 10 iterations. Sync all synchronizes the momentum of all layers.
Figure 9: Validation accuracy for training ResNet-18 on CIFAR100 and fine-tuning DeiT-Tiny on CIFAR100 using 8 workers on 2 nodes in our lower-bandwidth configuration. Interestingly, if you look at the training curve per iteration, even 1-bit allreduce closely mirrors Lion with 32-bit precision.
7 Discussion and Conclusion
Limitations
Distributed training methods, especially ones incorporating compression, are notoriously fickle, and their quality is often sensitive to the scale of the model and cluster being used. While we have demonstrated Lion Cub’s strong performance at a relatively small scale, computational resource limitations have so far prevented confirmation of its benefits at larger scales. We also note that we have focused on communication in a data-parallel setting, as opposed to model-parallelism techniques. While we believe Lion Cub could be extended to these approaches, we leave this to future work.
Another significant limitation is that while our method performs very well in lower-bandwidth networks, the improvements are less dramatic in low-latency, high-bandwidth networks. Despite this, we believe that distributed training in lower-bandwidth environments remains essential. We expect communication reduction to become increasingly important in the future. As deep learning becomes democratized, many more people will be training models on clusters with relatively low-end interconnects and will be communication-bound without improved techniques like Lion Cub. But even on high-end clusters, trends in hardware indicate that the balance of communication and computation will continue to worsen as GPU performance improves faster than network bandwidth, necessitating reduced communication for scalability and efficiency.
Conclusion
We presented Lion Cub, a suite of improvements enabling reduced communication overhead for distributed training with Lion. Lion Cub incorporates optimized communication primitives and a novel L1 quantization method, enabling it to converge comparably to full-precision Lion on a variety of models, including ResNets and GPT, while reducing overall runtime.
Acknowledgements
This work was performed under the auspices of the U.S. Department of Energy by Lawrence Livermore National Laboratory under Contract DE-AC52-07NA27344 (LLNL-TR-871246). T.B.N., B.V.E., and N.D. were funded by LLNL LDRD #24-SI-008. This work used computational resources TSUBAME4.0 supercomputer provided by Institute of Science Tokyo through the HPCI System Research Project (Project ID: hp240170) This work is supported by JST CREST Grant Number JPMJCR2112.
Appendix A Extended Related work
A.1 SignSGD
SignSGD is an optimization method that applies the $\operatorname{sign}$ operator to gradients, resulting in three possible values: +1, 0, and -1 (Seide et al., 2014; Bernstein et al., 2018, 2019). This approach has gained attention due to its similarity to Adam, particularly in terms of the normalized update (Kunstner et al., 2023), but also as a distributed learning method when combined with majority voting, as updates can be compressed (Bernstein et al., 2019). In majority voting, $\operatorname{sign}$ is first applied to the gradients, and then again to their sum: $\theta_{t+1}=\theta_{t}-\eta_{t}\operatorname{sign}\left[\sum_{i=1}^{N}\operatorname{sign}\nabla_{\theta}f_{i}(\theta_{t-1})\right].$ Note that $\operatorname{sign}$ is unbiased, which is why a stochastic $\operatorname{sign}$ is sometimes employed (Safaryan & Richtárik, 2021). Typically, discussions of majority voting assume a parameter server model for communication, where the signs computed on each worker are sent to a central server which computes the global sign and returns the result. As we will discuss, while straightforward, this approach does not scale.
A.2 Heavy-tailed noise and Gradient Clipping
We can regard our L1 quantization as performing quantization after gradient clipping. In heavy-tailed centralized training, gradient clipping is known to be an effective approach to mitigate heavy-tailed noise (Gorbunov et al., 2020; Zhang et al., 2020a). In distributed optimization, several works have shown that clipping gradients before communication helps curb the influence of heavy-tailed noise (Yang et al., 2022; Xiang & Su, 2023; Yan et al., 2024). FAT-Clipping integrates gradient clipping into federated averaging, either once per round or at each local step, to tame fat-tailed noise and guarantee convergence (Yang et al., 2022). Similarly, EPISODE applies episodic gradient clipping, which uses the global gradient norm at the start of each round to decide whether to clip local updates, to adaptively bound heavy-tailed stochastic noise and ensure robust convergence under data heterogeneity (Crawshaw et al., 2023). Furthermore, TQSGD first applies gradient clipping and then quantizes the clipped gradients for communication, thereby compensating for heavy-tailed noise, and is the method most closely related to our L1 quantization (Yan et al., 2024). In contrast to these methods, L1 quantization automatically performs both scaling and clipping without any additional hyperparameters, and further combines this mechanism with the Lion optimizer. Although we do not provide formal convergence guarantees, our empirical evaluation offers a practical demonstration of the importance of clipping before quantization in large-scale training.
A.3 Layerwise Selective Communication
We synchronize the momentum of selected layers every few iterations. Although no prior work has performed layer-wise selective momentum synchronization, federated learning has long cut communication costs by transmitting only certain layers’ weights or gradients more frequently or with higher precision. TWAFL is an asynchronous federated learning scheme in which the parameters of deep layers are updated less often than those of shallow layers (Chen et al., 2020). Likewise, FedLAMA reduces communication overhead by dynamically choosing each layer’s optimal aggregation interval by evaluating model discrepancy and communication costs (Lee et al., 2023). TLAQC focuses on quantizing and communicating only two representative layers of the model, while all other layers merely accumulate updates locally (Ren et al., 2023). In FedPart, each client selectively updates and transmits only a subset of the network (for example, just the feature‐extraction layers or only the classification layers) each round (Wang et al., 2024). Thus, in the federated‐learning, it has long been known that different layers exhibit different degrees of importance to synchronize. Although our approach keeps weights and gradients fully synchronized across all GPUs, letting only the momentum terms diverge, we have found that the amount of momentum synchronization required varies from layer to layer.
Appendix B Communication Design
B.1 The impact of 1-bit allreduces
In the second stage, an allgather disseminates aggregated updates, also with 1 bit. This is more challenging, because, in the typical case when there are an even number of workers, there can be an equal number of 1 and -1 updates for a parameter among the workers. Should this occur, the conversion to 1 bit can change the descent direction, impacting convergence. Standard Lion does not face this issue, as it applies the $\operatorname{sign}$ after summing updates. We analyzed the update vectors produced by standard Lion and the 1-bit allreduce when training a GPT model in Table 3. With 4 workers, over 30% of the parameters have an update where the majority vote is tied, leading to an erroneous update depending on how it is transmitted. While this rate decreases with 8 workers (as ties are less likely), it is still significant. Indeed, as Figure 10 demonstrates, these errors can lead to loss spikes and training divergence, and so are necessary to address, even if they may be relatively unlikely at scale. We find that a simple alternating sign method is sufficient to address this: we convert 0 to 1 on odd iterations and to -1 on even iterations before compressing to 1 bit. We considered a stochastic approach, but did not proceed with it as the overhead of random number generation was significant. Figure 10 shows that the alternating sign approach successfully stabilizes training and has minimal differences to using a 2-bit allreduce which transmits exact zeros.